OpenStudy (love_to_love_you):

Which hyperbola has both vertices in the same quadrant? Will fan and medal!

OpenStudy (love_to_love_you):

a) \[\frac{ (x+10)^2 }{11^2 }-\frac{ (y-3)^2 }{ 10^2 }=1\] b)\[\frac{ (y-8)^2 }{ 9^2}-\frac{ (x+6)^2 }{ 12^2 }=1\] c) \[\frac{ (x-1)^2 }{ 12^2}-\frac{ (y-11)^2 }{ 8^2 }=1\] d) \[\frac{ (y-10)^2 }{ 9^2 }-\frac{ (x+1)^2 }{ 11^2 }=1\]

OpenStudy (love_to_love_you):

@jigglypuff314

OpenStudy (love_to_love_you):

@robtobey

OpenStudy (love_to_love_you):

@SolomonZelman

OpenStudy (anonymous):

this problem is a pain in the neck! you have to find the vertices for each one!!

OpenStudy (anonymous):

if it was me, i would cheat vertices of the first one: \((-21,3),(1,3)\) so not that one

OpenStudy (anonymous):

answer is usually C, so i tried that next not that one either though